OpenStudy (anonymous):

subtract. simplify by collecting like radical terms if possible, assume that all number under radical represent non-negative number 3sqrt 81x - 3sqrt 3x^4

OpenStudy (anonymous):

\[3\sqrt{81x}-3\sqrt{3x^4}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[3\sqrt{81}\sqrt{x}-3\sqrt{x^4}\sqrt{3}\]

OpenStudy (anonymous):

\[3(9)\sqrt{x}-3(x^2)\sqrt{3}\]

OpenStudy (anonymous):

\[27\sqrt{x}-3x^2\sqrt{3}\]

OpenStudy (anonymous):

I'm pretty sure that's all you can do

OpenStudy (anonymous):

the 3 should be on top of the sqrt